# On the complexity of constrained reconfiguration and motion planning

### v1

@@ 1 Introduction
Going back to motion planning, we now introduce variants of problems on robotic arms under the umbrella of reconfiguration problems. Most previous results consider a small number of robots using a broad set of possible moves. We have chosen to restrict the power of robots to study the impact of the number and relative positions of robots on the complexity of the problem. We also focus on centralized algorithms, avoiding synchronization issues that might arise in multi-agent systems.
We introduce, among others, the Robotic Arm Motion Planning (RAMP) problem. The problem is intended to bridge the gap between reconfiguration problems that are geometric or have a natural planar representation, and reconfiguration problems on graphs. The RAMP problem is defined in the plane $\mathbb{R}^{2}$, with the canonical $x$-axis and $y$-axis, where a robotic arm is a unit-length line segment. An arm can be uniquely identified by its center, i.e., the midpoint $(x,y)$ of its line segment, and the angle $a\in[0,\pi)$ in radians that the line segment forms with the horizontal line passing through its center. A collection of arms is conflict-free whenever the corresponding line segments are pairwise non-intersecting. We are particularly interested in arms that are horizontal, which have an angle that is equal to 0, or vertical, which have an angle that is equal to $\pi/2$. A reconfiguration operation on a set of arms is a rotation of a single arm about its center, which corresponds to the robotic arm in question updating its angle $a$ to $a^{\prime}=(a+\alpha)\text{ mod }\pi$, where $\alpha\neq 0$. A rotation is said to be a counterclockwise rotation when $\alpha>0$, and is said to be a clockwise rotation otherwise. When a robotic arm rotates from one orientation to another, it sweeps two opposite sectors at the center of the arm, associated with the arcs defined between $a$ and $a^{\prime}$. If a robotic arm $r_{1}$ is rotating, and another arm $r_{2}$ intersects one of the two sectors swept by the rotation of $r_{1}$, we say that the rotation of $r_{1}$ hits $r_{2}$ (see Figure 1 for an illustration).
We now formally define the problems under consideration, starting with the most general problem:
@@
In order to utilize what is known in reconfiguration on graphs in terms of tractability and how it relates to a graph’s structure or its properties, it is reasonable to attempt to define both MM-RAMP and SM-RAMP in graph theoretic terms. Therefore, we consider two problems on graphs that generalize them, starting with the $k$-Compatible Ordering problem, which we define below:
@@
We show that $k$-Compatible Ordering has a range of applications that extends beyond our problems on robotic arms. The $k$-Compatible Ordering problem can be used to model problems that involve operations applied to elements from a reference set, where an operation either inserts an element into the set or removes one from it. For each element insertion or removal, we have choices for how the element should be inserted or removed, with constraints associated with each choice. As a concrete example, we are given a set of objects in containers, and we aim to transfer these objects into trucks in order to bring them to their final destinations. One possible way of doing so consists of completely emptying the containers and then filling up trucks. To save time and minimize the handling of heavy objects, we aim to load smaller trucks while unloading containers. However, there are constraints on how the trucks are to be filled. Trucks might have several destinations or certain objects might be fragile, resulting in constraints on the order of loading. Since trucks usually have several doors, there may be several possible ways to load the objects, which give rise to different constraints that can be represented by an instance of the $k$-Compatible Ordering problem as “$B$-type constraints”. Similarly, there are constraints on the unloading of the containers, which give rise to the “$A$-type constraints” in the $k$-Compatible Ordering problem.
It is easy to see that the SM-RAMP problem fits the criteria of the problems on containers described in the previous paragraph. We nonetheless provide a complete explanation for why SM-RAMP is a special case of the $k$-Compatible Ordering problem. Consider an instance of SM-RAMP. For each robotic arm $r$, we create a vertex $v$ in $V$. When we decide to rotate a robotic arm $r$, we can decide to rotate it either clockwise or counterclockwise. The constraints on the clockwise rotations will be represented by the pair $(A_{1},B_{1})$, while the constraints on the counterclockwise rotations will be represented by the pair $(A_{2},B_{2})$. More formally, if rotating robotic arm $r_{i}$ clockwise to its final orientation hits robotic arm $r_{j}$ in its initial orientation, then we add the arc $v_{i}v_{j}$ to the graph $A_{1}$. Similarly, if rotating robotic arm $r_{i}$ clockwise to its final orientation hits robotic arm $r_{j}$ in its final orientation, or if robotic arm $r_{j}$ in the final orientation intersects robotic arm $r_{i}$ in its initial orientation, then we add the arc $v_{j}v_{i}$ to $B_{1}$. The arcs in $(A_{2},B_{2})$ are defined similarly for the counterclockwise rotations (Figure 1).
@@
If the instance of SM-RAMP has a solution, then there is an ordering of the robotic arms and a labeling (clockwise or counterclockwise) of the robotic arms that satisfy all the constraints. Conversely, one can easily check that the ordering $\mathcal{S}$ and the labeling $\mathcal{L}$ of $2$-Compatible Ordering can be used to derive a solution to the SM-RAMP instance, with a clockwise rotation being associated with the pair $(A_{1},B_{1})$, and a counterclockwise rotation being associated with the pair $(A_{2},B_{2})$. Indeed, if a robotic arm $r_{i}$ represented by $s_{j}$ rotates in the clockwise direction, then its rotation should not hit any of the arms represented by $s_{1},\ldots,s_{j-1}$ in the vertical (final) orientation, which corresponds to $s_{j}$ being a source in $B_{1}[s_{1},\ldots,s_{j}]$, and its rotation should not hit any of the arms represented by $s_{j+1},\ldots,s_{n}$ which are still in their initial orientations, which corresponds to $s_{j}$ being a sink in $A_{1}[s_{j},\ldots,s_{n}]$.
SM-RAMP has a counterpart where robots can move back and forth before reaching their final orientations. One then wonders whether we can generalize MM-RAMP as we generalized SM-RAMP. The answer is positive. The following problem is a natural generalization of MM-RAMP:
@@
Unlike the $k$-Compatible Ordering problem, which allows only one move per agent (or robotic arm), the $k$-Compatible Set Arrangement problem allows multiple moves per agent and is therefore not surprisingly “harder” to solve. Our next result confirms this intuition by showing that the problem is indeed $\mathsf{PSPACE}$-complete.
@@ 3 Finding $k$-compatible orderings
Observations 123 and 4 greatly simplify the presentation of the hardness results in Section 3.2, as most of the gadgets that are used in the constructions depend on one of, or a combination of, these observations. For now, we have all the definitions needed to proceed to the positive results.
@@ 3.1 Positive Results
There are two directed graphs per label in the specification of the input to $k$-Compatible Ordering: one can wonder whether replacing each pair of directed graphs with the union of its graphs yields an equivalent instance (while maintaining the source and sink constraints on the resulting graphs). The answer is negative: Figure 2 justifies the use of two directed graphs per label in the input specification of $k$-Compatible Ordering. If we replace each graph pair with the union of the two graphs, and we treat all constraints as sink constraints (or, symmetrically, source constraints), the resulting instance has no valid labeled ordering, whereas a solution to the original instance is $\mathcal{S}=s_{1},s_{2},s_{3}$ and $\mathcal{L}=2,2,1$. In fact, substituting the $A$ graph and the $B$ graph of every pair with their union in the example of Figure 2 does not preserve solutions specifically because the union is a cycle. If this is not the case, we are dealing with a yes-instance of $k$-Compatible Ordering, and a valid labeled ordering can be computed efficiently.
Lemma 1.
Proof.
Let the labeling $\mathcal{L}$ be the labeling that assigns label $i$ to all vertices. We claim that we can find an ordering $\mathcal{S}$ of the vertices such that $(\mathcal{S},\mathcal{L})$ is a valid labeled ordering. We first compute a topological ordering for the union of $A_{i}$ and $B_{i}$ in time $\mathcal{O}(n+m)$ ($m$ being the number of edges in the union of $A_{i}$ and $B_{i}$). Let $\mathcal{S}=s_{1},\ldots,s_{n}$ be the topological ordering in reverse order. To show that $(\mathcal{S},\mathcal{L})$ is a solution, suppose, for the sake of contradiction, that there exists a vertex $s_{x}$ in $\mathcal{S}$ that is assigned label $i$ and that violates either the sink or the source constraint. If $s_{x}$ is not a sink in $A_{i}[s_{x},\ldots,s_{n}]$, then there exists an out-neighbor of $s_{x}$ in $s_{x+1},\ldots,s_{n}$. However, all the out-neighbors of $s_{x}$ in $A_{i}$ are in $s_{1},\ldots,s_{x-1}$, given that $\mathcal{S}$ is a reversed topological ordering. An analogous argument ensures that $s_{x}$ is a source in $B_{i}[s_{1},\ldots,s_{x}]$. Since the choice of $s_{x}$ is arbitrary, there is no vertex for which a source or sink constraint is violated, and $(\mathcal{S},\mathcal{L})$ is a valid labeled ordering, as needed. ∎
A polynomial algorithm for 1-Compatible Ordering can be derived almost immediately from the result of Lemma 1. In contrast to the example provided in Figure 2 where $k=2$, for 1-Compatible Ordering, an arc in $A_{1}$ and an arc in $B_{1}$ that share the same head and tail have the same interpretation, since both vertices will have the same label. In other words, whether an arc $uv$ exists in $A_{1}$, $B_{1}$, or both, it implies that, in a valid labeled ordering, $u$ occurs after $v$. This reduces the problem to finding a topological ordering on the union of $A_{1}$ and $B_{1}$.
@@
We prove that an instance $(V,\{(A_{1},B_{1})\})$ of 1-Compatible Ordering is a yes-instance if and only if the union of $A_{1}$ and $B_{1}$ is a directed acyclic graph. The backward direction has already been proved in Lemma 1.
As for the forward direction, towards a contradiction, suppose that $(V,\{(A_{1},B_{1})\})$ is a yes-instance, and the union of $A_{1}$ and $B_{1}$ contains at least one cycle. Consider an arbitrary solution $(\mathcal{S}=s_{1},\ldots,s_{n},\mathcal{L})$, where $\mathcal{L}$ trivially assigns label $1$ to all vertices. Since the union of $A_{1}$ and $B_{1}$ contains at least one cycle, there will exist a vertex $s_{x}$ in $\mathcal{S}$ that has an out-neighbor $s_{w}$ in $s_{x+1},\ldots,s_{n}$, otherwise, the union of $A_{1}$ and $B_{1}$ would have a topological ordering and would therefore be acyclic, a contradiction. Whether the arc $s_{x}s_{w}$ is in $A_{1}$ or $B_{1}$, it forces the solution to include $s_{w}$ before $s_{x}$, which contradicts $(\mathcal{S},\mathcal{L})$ being a solution to the instance.
Constructing $G^{\prime}$, the union of $A_{1}$ and $B_{1}$, can be done in time $\mathcal{O}(n+m)$ via a graph traversal of each of $A_{1}$ and $B_{1}$, and checking whether $G^{\prime}$ contains a cycle can be done via a depth-first search (DFS) in time $\mathcal{O}(n+m)$. If $G^{\prime}$ is acyclic, a topological ordering can be computed using two DFS calls. The running time of the algorithm is thus in $\mathcal{O}(n+m)$. ∎
The result from Lemma 2 can be generalized to obtain an algorithm that can be used to check whether a pair $(\mathcal{S}=s_{1},\ldots,s_{n},\mathcal{L}=\ell_{1},\ldots,\ell_{n})$ is a solution to an instance of $k$-Compatible Ordering in time linear in the number of vertices and the number of constraint edges incident to the $s_{i}$’s in the directed graphs specified by their label, i.e., only the outgoing edges of $s_{i}$ in $A_{\ell_{i}}$ and the incoming edges of $s_{i}$ in $B_{\ell_{i}}$ are accounted for. This algorithm, which is directly implied by Observation 5 below and involves checking for acyclicity via topologically ordering the vertices, is asymptotically optimal, and will be used as a subroutine in Theorems 4 and 5.
Observation 5.
@@
Given that Lemmas 1 and 2 make use of the acyclicity of the union of a specific pair of graphs $(A_{i},B_{i})$ in $\mathcal{G}$ to detect yes-instances of $k$-Compatible Ordering, it is natural to consider whether we can hope to solve instances that include cycles, whether in a given directed graph, in the union of two directed graphs $(A_{i},B_{i})$, or in the union of $\mathcal{G}$, in polynomial time. The answer to those questions is negative, and the relevant results are presented in Section 3.2.
@@
    Case 4: The compcomp values of an introduce node $N_{i}$ with child $N_{j}$ can be computed in time $\mathcal{O}(t^{2}\cdot t!\cdot k^{t})$. We go through each of the $\mathcal{O}(t!\cdot k^{t})$ labeled orderings $(\mathcal{S}_{i},\mathcal{L}_{i})$ of $V(N_{i})$ and we remove $v$ to obtain the labeled ordering $(\mathcal{S}_{i}-v,\mathcal{L}_{i}-v)$ in time $\mathcal{O}(t)$. If $comp(N_{j},\mathcal{S}_{i}-v,\mathcal{L}_{i}-v)=0$ or if $(\mathcal{S}_{i},\mathcal{L}_{i})$ is not a solution to $\mathcal{I}(N_{i})$, we set $comp(N_{i},\mathcal{S}_{i},\mathcal{L}_{i})=0$; otherwise, we set $comp(N_{i},\mathcal{S}_{i},\mathcal{L}_{i})=1$. Both the construction of $\mathcal{I}(N_{i})$ and verifying whether $(\mathcal{S}_{i},\mathcal{L}_{i})$ is a solution can be done in time linear in its size, i.e., in time $\mathcal{O}(t^{2})$ (Observation 5), since $\mathcal{I}(N_{i})$ has $\mathcal{O}(t)$ vertices and $\mathcal{O}(t^{2})$ arcs.
@@
    As such, we set $V_{1}=V_{T}(N_{g}),V_{2}=V_{T}(N_{h})$. By Observation 7, all the preconditions of Observation 6 hold; we can therefore apply Observation 6 to prove the correctness of the compcomp values of $N_{i}$ that are assigned the value 1.
@@
Suppose that $G^{f}$ is obtained via an application of O4, by substituting the labeled directed graphs $G^{f}_{1},\ldots,G^{f}_{p}$ into the template graph $G_{t}$ that has vertex set $v_{1},\ldots,v_{p}$. To compute which label selections can be used to solve $G^{f}$, we look at the template graph, we go over all possible label selections of the $p$ instances, and we check which combinations of label selections allow for an ordering of the vertices of $G_{t}$. Formally, for each element $(L_{1},\ldots,L_{p})$ in $\mathbb{L}(G^{f}_{1})\times\ldots\times\mathbb{L}(G^{f}_{p})$, we update the neighborhood of each vertex $v_{j}$ in $V(G_{t})$ to preserve $v_{j}$’s outgoing and incoming arcs whose label set shares at least one of the labels in $L_{j}$, and delete the rest of the arcs. Then, we check whether $G_{t}$ is a directed acyclic graph. If it is, $L_{1}\cup\ldots\cup L_{p}$ is a label selection that can be used to solve the instance induced on the vertices of $G^{f}$; if it is not, then it is not possible to solve the instance induced on the vertices of $G^{f}$ using only labels from $L_{1}\cup\ldots\cup L_{p}$ (Observation 5).

### v3

@@ 1 Introduction
Going back to motion planning, we now present variants of problems on robotic arms under the umbrella of reconfiguration problems. Most previous results consider a small number of robots using a broad set of possible moves. We have chosen to restrict the power of robots to study the impact of the number and relative positions of robots on the complexity of the problem. We also focus on centralized algorithms, avoiding synchronization issues that might arise in multi-agent systems. We refer to the most general setup as the Robotic Arm Motion Planning (RAMP) problem.
The presentation of the RAMP problem, followed by its variants, is meant to provide context for and a motivation of the paper’s two main problems, namely the $k$-Compatible Ordering problem and the $k$-Compatible Set Arrangement problem, with the included results for these two problems being used to refine what we know about the tractability of different versions of the RAMP problem. RAMP problems will therefore be recurrently used as tangible examples of $k$-compatible ordering and arrangement problems.
The RAMP problem is defined in the plane $\mathbb{R}^{2}$, with the canonical $x$-axis and $y$-axis, where a robotic arm is a unit-length line segment. An arm can be uniquely identified by its center, i.e., the midpoint $(x,y)$ of its line segment, and the angle $a\in[0,\pi)$ in radians that the line segment forms with the horizontal line passing through its center. A collection of arms is conflict-free whenever the corresponding line segments are pairwise non-intersecting. We are particularly interested in arms that are horizontal, which have an angle that is equal to 0, or vertical, which have an angle that is equal to $\pi/2$. A reconfiguration operation on a set of arms is a rotation of a single arm about its center, which corresponds to the robotic arm in question updating its angle $a$ to $a^{\prime}=(a+\alpha)\text{ mod }\pi$, where $\alpha\neq 0$. A rotation is said to be a counterclockwise rotation when $\alpha>0$, and is said to be a clockwise rotation otherwise. When a robotic arm rotates from one orientation to another, it sweeps two opposite sectors at the center of the arm, associated with the arcs defined between $a$ and $a^{\prime}$. If a robotic arm $r_{1}$ is rotating, and another arm $r_{2}$ intersects one of the two sectors swept by the rotation of $r_{1}$, we say that the rotation of $r_{1}$ hits $r_{2}$ (see Figure 1 for an illustration).
We now formally define the problems under consideration, starting with the most general formulation of the RAMP problem:
@@
In order to utilize what is known in reconfiguration on graphs in terms of tractability and how it relates to a graph’s structure or its properties, it is reasonable to attempt to define both MM-RAMP and SM-RAMP in graph theoretic terms. Therefore, we consider two problems on graphs that generalize them, starting with the $k$-Compatible Ordering problem. The $k$-Compatible Ordering problem, which is a generalization of SM-RAMP, is defined below:
@@
Observation 1.
There exists a polynomial-time reduction from the SM-RAMP problem to the $2$-Compatible Ordering problem.
Proof.
Consider an instance $\mathcal{R}=r_{1},\ldots,r_{n}$ of SM-RAMP. For each robotic arm $r_{i}\in\mathcal{R}$, we create a vertex $v_{i}$ in $V$. If the clockwise (resp. counterclockwise) rotation of an arm $r_{i}$ hits an arm $r_{j}$, we add the arc $r_{i}r_{j}$ to $A_{1}$ (resp. $A_{2}$). If the vertical orientation of an arm $r_{i}$ is hit by the clockwise (resp. counterclockwise) rotation of an arm $r_{j}$, or crosses $r_{j}$ in its initial orientation, we add the arc $r_{i}r_{j}$ to $B_{1}$ (resp. $B_{2}$). The vertex set $V$ and the four directed graphs $A_{1}$, $B_{1}$, $A_{2}$, $B_{2}$ can clearly be constructed in time polynomial in $n$, the number of robotic arms in $\mathcal{R}$. We claim that $\mathcal{R}$ is a yes-instance of SM-RAMP if and only if $\{(V,\mathcal{G}=\{(A_{1},B_{1}),(A_{2},B_{2})\}$ is a yes-instance of $2$-Compatible Ordering. We only present the forward direction of the proof, as the argument for the backward direction is symmetrical and follows directly from the construction of the $2$-Compatible Ordering instance.
Let $(\mathcal{R}_{\text{sol}},\mathcal{D})$ represent a solution to the SM-RAMP instance. More specifically, let $\mathcal{R}_{\text{sol}}=r^{\prime}_{1},\ldots,r^{\prime}_{n}$ be an ordering of the robotic arms in $\mathcal{R}$ such that $r^{\prime}_{i}$ is the $i$th robotic arm to be rotated, for $1\leq i\leq n$, and let $\mathcal{D}=d_{1},\ldots,d_{n}$ denote the direction in which the robotic arms are rotated, such that $d_{i}\in\{r,l\}$, for $1\leq i\leq n$ (where $r$ and $l$ denote a clockwise and a counterclockwise rotation, respectively). We claim that $(\mathcal{S}=v^{\prime}_{1},\ldots,v^{\prime}_{n},\mathcal{L}=\ell_{1},\ldots,\ell_{n})$, where $v^{\prime}_{i}$ is the vertex associated with $r^{\prime}_{i}$ in the construction, and where $\ell_{i}=1$ if $d_{i}=r$ and $\ell_{i}=2$ if $d_{i}=l$, is a valid labeled ordering for the constructed SM-RAMP instance. Assume otherwise, then, there must exist two vertices $v^{\prime}_{i}$ and $v^{\prime}_{j}$ in $\mathcal{S}$ (without loss of generality, $i<j$) that are assigned labels $\ell_{i}$ and $\ell_{j}$ such that either $v^{\prime}_{i}v^{\prime}_{j}$ is an arc in $A_{\ell_{i}}$, or $v^{\prime}_{i}v^{\prime}_{j}$ is an arc in $B_{\ell_{j}}$. However, by construction, the existence of the former of these two arcs implies that the rotation of $r^{\prime}_{i}$ in direction $d_{i}$ would have hit $r^{\prime}_{j}$ in its initial orientation, and that of the latter of these two arcs implies that the rotation of $r^{\prime}_{j}$ in direction $d_{j}$ would have hit $r^{\prime}_{i}$ in its vertical orientation, or that $r^{\prime}_{i}$ in its vertical orientation crosses $r^{\prime}_{j}$ in its initial orientation: either of these two cases invalidates $(\mathcal{R}_{\text{sol}},\mathcal{D})$ being a solution to the SM-RAMP instance. ∎
While the SM-RAMP problem, which we initially set to solve, inspired the $k$-Compatible Ordering problem, the latter’s applications extend beyond problems on robotic arms, as it can capture constraints that are not inherently geometric. More precisely, the $k$-Compatible Ordering problem models a class of element-ordering problems in which each element admits one of $k$ possible insertion modes, where the chosen mode determines the set of elements that must precede it and the set that must follow it, independently of the insertion modes of these other elements.
We use the Truck Unloading problem as an example of a $3$-Compatible Ordering problem. Let $b_{1},\ldots,b_{n}$ be a set of boxes in a truck, where each box is a rectangular cuboid. We are looking to place these boxes in a single stack on a platform inside a warehouse, while ensuring that the stack does not topple. Each box may be placed in the stack with any of its six faces oriented downward; we will assume that opposite faces count as the same orientation, so each box has three orientations, each of which corresponds to a distinct face oriented downward (where no two of those distinct faces are opposite to each other). The three faces are numbered $1$ to $3$ arbitrarily. The orientation of a box determines its weight distribution, which should be taken into account when it comes to balancing the stack. Choosing label $\ell$ for box $b_{i}$ means: “put box $b_{i}$ in the stack with face $\ell$ facing downward”. Getting a box onto the stack on the platform is a two-step process: the first step involves getting the box out of the truck, and the second step involves placing the box at the top of the stack, in such a way that does not cause the stack to topple. We add one more restriction: the orientation of the box has to be chosen before it is lifted (i.e., before the first step), and the box preserves its orientation until it is placed on top of the stack. $A$ graph constraints are associated with the first step, whereas $B$ graph constraints are associated with the second step. More specifically, the graph $A_{\ell}$ has an arc $b_{i}b_{j}$ whenever moving $b_{i}$ outside of the truck in orientation $\ell$ must wait until $b_{j}$ has already been moved onto the stack (e.g., $b_{i}$ sits deeper in the truck than $b_{j}$, and cannot be carried out of the truck in orientation $\ell$ without hitting $b_{j}$). The graph $B_{\ell}$ has an arc $b_{i}b_{j}$ whenever $b_{i}$ being added to the stack in any orientation before $b_{j}$ in orientation $\ell$ would cause the stack to topple, due to the weight distribution in $b_{i}$ prohibiting $b_{j}$ from being placed on top of it in orientation $\ell$, irrespective of $b_{i}$’s orientation. Thus, when a box $b_{i}$ is assigned label $\ell$ in a solution’s ordering, we simultaneously enforce that all its $A_{\ell}$-successors appear earlier than it in the ordering, ensuring that no box prevents box $b_{i}$ from being moved out of the truck, and all its $B_{\ell}$-predecessors appear later than it in the ordering of $V$ in a solution, ensuring that no box positioned under $b_{i}$ in the stack causes the stack to topple.
While both the SM-RAMP problem and the Truck Unloading problem can be modeled as $k$-Compatible Ordering problems ($k=2$ for SM-RAMP, $k=3$ for Truck Unloading), the value of $k$ is fixed and is intrinsic to the definition of both problems. With that being said, Truck Unloading can be generalized more naturally by replacing boxes with polyhedrons that can be oriented in one of $k$ ways. We have therefore decided to include the parameter $k$ in the formulation of $k$-Compatible Ordering not only to capture a wider array of problems, but also to analyze how the value of $k$ affects the problem’s tractability.
@@
SM-RAMP has a counterpart where robots can move back and forth before reaching their final orientations, which is the MM-RAMP problem defined earlier. One then wonders whether we can generalize MM-RAMP as we generalized SM-RAMP. The answer is positive. The following problem is a natural generalization of MM-RAMP:
@@
In an attempt to narrow down what makes $k$-Compatible Ordering hard, we consider whether looking for an orderable subset of the vertices, subject to the same constraints, makes the problem any easier to solve. In particular, we investigate whether the problem can be solved in polynomial time when the subset is bounded in size. We prove that this is likely false.
Theorem 6.
Our final result concerns the $k$-Compatible Set Arrangement problem. Unlike the $k$-Compatible Ordering problem, which allows only one move per agent (or robotic arm), the $k$-Compatible Set Arrangement problem allows multiple moves per agent and is therefore not surprisingly “harder” to solve. Our next result confirms this intuition by showing that the problem is indeed $\mathsf{PSPACE}$-complete.
@@ 3 Finding $k$-compatible orderings
Observations 234 and 5 greatly simplify the presentation of the hardness results in Section 3.2, as most of the gadgets that are used in the constructions depend on one of, or a combination of, these observations. For now, we have all the definitions needed to proceed to the positive results.
@@ 3.1 Positive Results
There are two directed graphs per label in the specification of the input to $k$-Compatible Ordering: one can wonder whether replacing each pair of directed graphs with the union of its graphs yields an equivalent instance (while maintaining the source and sink constraints on the resulting graphs). The answer is negative: Figure 2 justifies the use of two directed graphs per label in the input specification of $k$-Compatible Ordering. If we replace each graph pair with the union of the two graphs, and we treat all constraints as sink constraints (or, symmetrically, source constraints), the resulting instance has no valid labeled ordering, whereas a solution to the original instance is $\mathcal{S}=s_{1},s_{2},s_{3}$ and $\mathcal{L}=2,2,1$. In fact, substituting the $A$ graph and the $B$ graph of every pair with their union in the example of Figure 2 does not preserve solutions specifically because the union is a cycle. If this is not the case, we are dealing with a yes-instance of $k$-Compatible Ordering, and a valid labeled ordering can be computed efficiently (Corollary 1). We first show how checking for cycles in the union of the graphs is sufficient to determine whether a $1$-Compatible Ordering instance is a yes-instance.
@@
We prove that an instance $(V,\{(A_{1},B_{1})\})$ of 1-Compatible Ordering is a yes-instance if and only if the union of $A_{1}$ and $B_{1}$ is a directed acyclic graph.
For the forward direction, towards a contradiction, suppose that $(V,\{(A_{1},B_{1})\})$ is a yes-instance, and that the union of $A_{1}$ and $B_{1}$ contains at least one cycle. Consider an arbitrary solution $(\mathcal{S}=s_{1},\ldots,s_{n},\mathcal{L})$, where $\mathcal{L}$ trivially assigns label $1$ to all vertices. Since the union of $A_{1}$ and $B_{1}$ contains at least one cycle, there will exist a vertex $s_{x}$ in $\mathcal{S}$ that has an out-neighbor $s_{w}$ in $s_{x+1},\ldots,s_{n}$, otherwise, the union of $A_{1}$ and $B_{1}$ would have a topological ordering and would therefore be acyclic, a contradiction. Whether the arc $s_{x}s_{w}$ is in $A_{1}$ or $B_{1}$, it forces the solution to include $s_{w}$ before $s_{x}$, which contradicts $(\mathcal{S},\mathcal{L})$ being a solution to the instance.
As for the backward direction, let the labeling $\mathcal{L}$ be the labeling that assigns label $1$ to all vertices. We claim that we can find an ordering $\mathcal{S}$ of the vertices such that $(\mathcal{S},\mathcal{L})$ is a valid labeled ordering. We first compute a topological ordering for the union of $A_{1}$ and $B_{1}$ in time $\mathcal{O}(n+m)$ ($m$ being the number of edges in the union of $A_{1}$ and $B_{1}$). Let $\mathcal{S}=s_{1},\ldots,s_{n}$ be the topological ordering in reverse order. To show that $(\mathcal{S},\mathcal{L})$ is a solution, suppose, for the sake of contradiction, that there exists a vertex $s_{x}$ in $\mathcal{S}$ that is assigned label $1$ and that violates either the sink or the source constraint. If $s_{x}$ is not a sink in $A_{1}[s_{x},\ldots,s_{n}]$, then there exists an out-neighbor of $s_{x}$ in $s_{x+1},\ldots,s_{n}$. However, all the out-neighbors of $s_{x}$ in $A_{i}$ are in $s_{1},\ldots,s_{x-1}$, given that $\mathcal{S}$ is a reversed topological ordering. An analogous argument ensures that $s_{x}$ is a source in $B_{1}[s_{1},\ldots,s_{x}]$. Since the choice of $s_{x}$ is arbitrary, there is no vertex for which a source or sink constraint is violated, and $(\mathcal{S},\mathcal{L})$ is a valid labeled ordering, as needed. ∎
A generalization of the backward direction of Lemma 1 can be used to characterize trivial yes-instances of $k$-Compatible Ordering as follows:
Corollary 1.
The result from Corollary 1 can also be used to obtain an asymptotically optimal algorithm that can be used to check whether a pair $(\mathcal{S}=s_{1},\ldots,s_{n},\mathcal{L}=\ell_{1},\ldots,\ell_{n})$ is a solution to an instance of $k$-Compatible Ordering in time linear in the number of vertices and the number of constraint edges incident to the $s_{i}$’s in the directed graphs specified by their label, i.e., only the outgoing edges of $s_{i}$ in $A_{\ell_{i}}$ and the incoming edges of $s_{i}$ in $B_{\ell_{i}}$ are accounted for. This algorithm will be used as a subroutine in Theorems 4 and 5.
Observation 6.
@@
Given that Corollary 1 makes use of the acyclicity of the union of a specific pair of graphs $(A_{i},B_{i})$ in $\mathcal{G}$ to detect yes-instances of $k$-Compatible Ordering, it is natural to consider whether we can hope to solve instances that include cycles, whether in a given directed graph, in the union of two directed graphs $(A_{i},B_{i})$, or in the union of $\mathcal{G}$, in polynomial time. The answer to those questions is negative, and the relevant results are presented in Section 3.2.
@@
    Case 4: The compcomp values of an introduce node $N_{i}$ with child $N_{j}$ can be computed in time $\mathcal{O}(t^{2}\cdot t!\cdot k^{t})$. We go through each of the $\mathcal{O}(t!\cdot k^{t})$ labeled orderings $(\mathcal{S}_{i},\mathcal{L}_{i})$ of $V(N_{i})$ and we remove $v$ to obtain the labeled ordering $(\mathcal{S}_{i}-v,\mathcal{L}_{i}-v)$ in time $\mathcal{O}(t)$. If $comp(N_{j},\mathcal{S}_{i}-v,\mathcal{L}_{i}-v)=0$ or if $(\mathcal{S}_{i},\mathcal{L}_{i})$ is not a solution to $\mathcal{I}(N_{i})$, we set $comp(N_{i},\mathcal{S}_{i},\mathcal{L}_{i})=0$; otherwise, we set $comp(N_{i},\mathcal{S}_{i},\mathcal{L}_{i})=1$. Both the construction of $\mathcal{I}(N_{i})$ and verifying whether $(\mathcal{S}_{i},\mathcal{L}_{i})$ is a solution can be done in time linear in its size, i.e., in time $\mathcal{O}(t^{2})$ (Observation 6), since $\mathcal{I}(N_{i})$ has $\mathcal{O}(t)$ vertices and $\mathcal{O}(t^{2})$ arcs.
@@
    As such, we set $V_{1}=V_{T}(N_{g}),V_{2}=V_{T}(N_{h})$. By Observation 8, all the preconditions of Observation 7 hold; we can therefore apply Observation 7 to prove the correctness of the compcomp values of $N_{i}$ that are assigned the value 1.
@@
Suppose that $G^{f}$ is obtained via an application of O4, by substituting the labeled directed graphs $G^{f}_{1},\ldots,G^{f}_{p}$ into the template graph $G_{t}$ that has vertex set $v_{1},\ldots,v_{p}$. To compute which label selections can be used to solve $G^{f}$, we look at the template graph, we go over all possible label selections of the $p$ instances, and we check which combinations of label selections allow for an ordering of the vertices of $G_{t}$. Formally, for each element $(L_{1},\ldots,L_{p})$ in $\mathbb{L}(G^{f}_{1})\times\ldots\times\mathbb{L}(G^{f}_{p})$, we update the neighborhood of each vertex $v_{j}$ in $V(G_{t})$ to preserve $v_{j}$’s outgoing and incoming arcs whose label set shares at least one of the labels in $L_{j}$, and delete the rest of the arcs. Then, we check whether $G_{t}$ is a directed acyclic graph. If it is, $L_{1}\cup\ldots\cup L_{p}$ is a label selection that can be used to solve the instance induced on the vertices of $G^{f}$; if it is not, then it is not possible to solve the instance induced on the vertices of $G^{f}$ using only labels from $L_{1}\cup\ldots\cup L_{p}$ (Observation 6).